OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? x^4 – 41x^2 = –400 I need step by step help.

OpenStudy (anonymous):

I know that I have to get -400 on the other side and by doing so it would be x^4-41x^2+400=0 and then I let a=x^2 making a^2-41a+400=0 and now I have to factor which confuses me